# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v2

@@ Abstract
There has been a growing interest in studying sequential neural posterior estimation (SNPE) techniques for their advantages in dealing with simulation-based models with intractable likelihoods. They are devoted to learning the posterior from adaptively proposed simulations using neural network-based conditional density estimators. As a SNPE technique, the automatic posterior transformation (APT) method proposed by Greenberg et al. (2019) performs notably and scales to high dimensional data. However, the APT method bears the computation of an expectation of the logarithm of an intractable normalizing constant, i.e., a nested expectation. Although atomic APT was proposed to solve this by discretizing the normalizing constant, it remains challenging to analyze the convergence of learning. In this paper, we propose a nested APT method to estimate the involved nested expectation instead. This facilitates establishing the convergence analysis. Since the nested estimators for the loss function and its gradient are biased, we make use of unbiased multi-level Monte Carlo (MLMC) estimators for debiasing. To further reduce the excessive variance of the unbiased estimators, this paper also develops some truncated MLMC estimators by taking account of the trade-off between the bias and the average cost. Numerical experiments for approximating complex posteriors with multimodal in moderate dimensions are provided.
@@ 1 Introduction
Simulator-based models are widely used across various scientific disciplines, including neuroscience [46], physics [7, 23], biology [28, 38, 45] , and inverse graphics [52]. These models serve as crucial tools for describing and comprehending investigated processes based on observed data. However, when applying traditional Bayesian inference to simulator-based models, challenges arise, such as the intractable likelihood function $p(x|\theta)$ and the computational expense associated with running the simulator.
@@
To address these challenges, a series of likelihood-free Bayesian computation (LFBC) methods have been developed. These methods include approximate Bayesian computation (ABC) [5, 44], synthetic likelihoods (SL) [50, 58], Bayes optimization [26], likelihood-free inference by ratio estimation [56], and pseudo marginals methods [2, 3]. A comprehensive summary and review of these methods can be found in [10], and they have all been benchmarked in [31, 40].
@@
Posterior density estimation approaches approximate the posterior of interest $p(\theta|x_{o})$ with a family of density estimators $q_{\phi}(\theta)$, where $\phi$ is the parameters of the density estimators. Optimization-based approaches are widely used in these methods, and the Kullback-Leibler (KL) divergence between $p(\theta|x)$ and $q_{\phi}(\theta)$, which measures the differences between two densities, is commonly chosen as the loss function. Variational Bayes (VB), as a computationally optimization-based effective method for approximating the posterior distribution of a Bayesian problem, is widely used. In the likelihood-free context, [57] developed a new VB method with an intractable likelihood, while [29] proposed an unbiased VB method based on nested MLMC. However, these methods tend to fail in cases where simulations are expensive, as the nested estimation in these methods requires additional simulation procedures.
@@
For this reason, there has been a growing interest in employing neural networks to represent probability density recently, particularly normalizing flows [37]. When the inference problem is focused solely on the observation $x_{o}$, the data efficiency can be improved using the sequential training schemes proposed in the sequential neural posterior estimation (SNPE) methods [24, 41, 47]. In this approach, model parameters are drawn from a proposal distribution that is more informative about $x_{o}$ compared to the prior distribution. However, SNPE requires a modification of the loss function compared to neural posterior estimation (NPE) to ensure that the neural network approximates the true posterior $p(\theta|x)$. Many approaches have been developed to address this issue.
Among the SNPE methods, automatic posterior transformation (APT) requires the computation of the expectation of the logarithm of an intractable normalizing constant, which is a nested expectation. They then propose to use the atomic APT to discretize the normalizing constant. However, it remains challenging to utilize the existing analysis techniques [6] to perform convergence analysis to the best of our knowledge. To address this limitation, we propose the nested APT method, a method enable the convergence analysis without sacrificing the performance. We conduct a thorough analysis of the bias and variance of the loss estimators, moreover, we provide a convergence result for this biased estimator in the case of stochastic gradient descent (SGD). To evaluate the effectiveness of the nested APT, we conduct a series of comprehensive numerical experiments.
Since nested APT is a biased estimator, we undertake a comprehensive exploration of a range of unbiased methods based on multi-level Monte Carlo (MLMC) [16] to eliminate the bias. We establish several theorems that investigate the order of bias, variance, and the average cost of the estimators of the loss function and its gradient. Additionally, as we use stochastic gradient-based optimization methods to update the parameters of the neural network, it is well known that the unbiased method suffers from excessive variance in the gradient. To address this issue, we employ several variants of truncated MLMC methods to diminish the variance. Furthermore, we compute the order of the variance and average cost for the estimators for the gradient of the loss function. Finally, we utilize these results to provide the convergence analysis of truncated MLMC and nested APT.
@@
The remainder of this paper is organized as follows. In Section 2, we propose the nested APT method and provide insights into possible failures when directly using the nested APT. Moreover, we propose a strategy to address this failure. We also present the experimental results and analyze them in this section. In Section 3, we introduce the basics of MLMC methods, including their formulation and theoretical analysis of the random sequence utilized in MLMC methods. Equipped with these tools, we develop both unbiased MLMC methods and truncated MLMC methods and analyze the order of variance for the losses and gradients, as well as the average cost. In Section 4, we provide convergence analysis for nested APT and truncated MLMC methods in the case of SGD. In  Section 5, we conduct a series of numerical experiments on benchmark tasks. The results of these experiments are presented and analyzed. Finally, we conclude this paper with some remarks in Section 6. A lengthy proof and some numerical results are deferred in the appendices.
@@ 2.1 Problem formulation
Assume that the prior distribution has a density $p(\theta)$ with respect to the Lebesgue measure, where the model parameter of interest $\theta\in\Theta$. We focus on the finite-dimensional parameter space $\Theta$. Given an observed sample $x_{o}$, our objective is to perform inference on the posterior $p(\theta|x_{o})\propto p(\theta)p(x_{o}|\theta)$. However, in many cases, the likelihood $p(x|\theta)$ either lacks an explicit expression or presents challenges in direct evaluation. Instead, it can be expressed using a ‘simulator’ approach, wherein, given a fixed model parameter $\theta$, we can sample $x$ from $p(x|\theta)$.
@@
which is not less than 00 by using Jensen’s inequality and attains its minima when $q(\theta)$ agrees with $p(\theta)$, making it suitable as a loss function. When the likelihood $p(x|\theta)$ is tractable, one can directly approximate $p(\theta|x_{o})$ by minimizing the KL divergence between the target distribution and the proposed estimator $q(\theta)$ within a certain family of distributions. In the likelihood-free context, this can be viewed as a problem of conditional density estimation. Within this framework, a conditional density estimator $q_{F(x,\phi)}(\theta)$ based on a neural network [47, 24], is utilized to approximate $p(\theta|x)$ over the admissible set of tuning parameter $\phi\in\Phi$, where $\Phi\subset\mathbb{R}^{d}$ is the space of the neural network parameter with a finite dimension $d$. To this end, we focus on minimizing the following average KL divergence under the marginal distribution with density $p(x)=\int p(\theta)p(x|\theta)\mathrm{d}\theta$
@@
is used as the loss function. However, since $L_{1}(\phi)$ is intractable, we use its empirical estimator instead
@@
Since we aim to conduct conditional density estimation at $x_{o}$, it is essential to utilize a proposal $\tilde{p}(\theta)$ that provides more informative priors regarding $x_{o}$ in comparison to the prior $p(\theta)$. After initializing $\tilde{p}(\theta)$ as $p(\theta)$, we then want the approximation of $p(\theta|x_{o})$ to serve as a good candidate of the proposal in the following simulation. This conditional density estimation with an adaptively chosen proposal is called sequential neural posterior estimation (SNPE). However, after replacing $p(\theta,x)$ with $\tilde{p}(\theta,x)=\tilde{p}(\theta)p(x|\theta)$ in (1), it is observed that $q_{F(x,\phi)}(\theta)$ approximates the proposal posterior
@@
where $\tilde{p}(x)=\int\tilde{p}(\theta)p(x|\theta)\mathrm{d}\theta$. Hence, we need to adjust the loss function to make $q_{F(x,\phi)}(\theta)$ approximate the true posterior $p(\theta|x)$.
@@
Since the integral $Z(x,\phi)$ is usually intractable in practice, APT proposes the use of ‘atomic’ proposals. APT with such proposals is known as atomic APT. Specifically, they assume that $\tilde{p}(\theta)$ is a discrete uniform distribution $\mathcal{U}\{\theta_{1},\dots,\theta_{M}\}$, where $\theta_{i}$ are sampled from some distribution. The uniform setting of the proposal distribution allows for the analytical computation of $Z(x,\phi)$. With this method, (3) and (5) can be reformulated as
@@
However, to the best of our knowledge, the use of atomic proposals makes it challenging to analyze their convergence behavior. Therefore we are unable to explain the unexpected low performance in some tasks [12] with the existing convergence results [6], both in terms of insight and theory. As an alternative approach to estimating $Z(x,\phi)$, the nested estimation, which enjoys a comprehensive theoretical framework and comparable performance, is studied in the next section.
@@ 2.2 Nested estimation
is selected as the loss function, and the last term $-\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$ in (7) is dropped for it is independent of $\phi$. We employ stochastic gradient methods to optimize the loss function in this paper, and therefore $\nabla_{\phi}\mathcal{L}(\phi)$ is of interest. By allowing the interchange of expectation $\mathbb{E}[\cdot]$ and gradient operator $\nabla_{\phi}$, the gradient of (8) is given by
@@
Notice that (9) holds a similar formulation of the gradient used in the optimization procedure of Bayesian experimental design [9, 15, 33, 36]. Given any query point $(\theta,x)\sim\tilde{p}(\theta,x)$, we denote the corresponding query of the loss and gradient as
@@
Given the intractability of the normalizing constant $Z(x,\phi)=\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]$, it remains an obstacle to derive an estimator for the loss function. A simple choice is leveraging its empirical estimator based on $M$ samples, which is given by
@@
^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ $=\frac{1}{N}\sum_{i=1}^{N}\psi_{\phi,M}^{(i)}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{M}\sum_{j=1}^{M}\frac{g_{\phi}(x_{i},\theta^{\prime}_{ij})}{g_{\phi}(x_{i},\theta_{i})},$ (15)
@@
where $\theta_{ij}^{\prime}\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta^{\prime})$ constitute the inner sample with a size of $M$ and $(\theta_{i},x_{i})\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta,x)$ constitute the outer sample with a size of $N$, $\psi^{(i)}_{\phi,M},\ \rho^{(i)}_{\phi,M}$ are iid copies of $\psi_{\phi,M}$ and $\rho_{\phi,M}$ respectively. Due to the nonlinearity of the logarithm, (15) is a biased estimator. Using similar arguments in [53], it is not difficult to evaluate that the nested estimator ^$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ has a variance of $\mathcal{O}(1/N)$ and a bias of $\mathcal{O}(1/M)$. For a given mean squared error (MSE) $\epsilon$, choosing $N=\mathcal{O}(\epsilon^{-2})$ and $M=\mathcal{O}(\epsilon^{-1})$ yields a computational complexity of $\mathcal{O}(NM)=\mathcal{O}(\epsilon^{-3})$.
Compared to nested APT, the bias properties of atomic APT are less straightforward to analyze theoretically. However, empirical results show that the performance of both methods is similar. Visual comparisons of their approximated posteriors are provided in Appendix C. However, since nested APT updates the model with a biased gradient, the update direction will systematically deviate from the optimal solution of the loss function, potentially preventing the model from converging to the global optimum or leading it to an incorrect local extremum, increasing training time and computational cost. What’s worse, the parameter updates may fail to reflect the characteristics of the target posterior. To address this issue, we propose to make use of the existing MLMC techniques for the improvement of the nested estimator.
@@ 3.1 Basic idea of MLMC
We first introduce the idea of MLMC in a generic setting following [16]. Consider the problem of estimating $\mathbb{E}[P]$, where the sampling of $P$ is costly or even infeasible. Suppose that $P$ can be approximated by tractable $P_{\ell}$ with increasing accuracy, but also increasing cost, as $\ell\to\infty$. As a result, the expectation $\mathbb{E}[P]$ can be estimated by the simple Monte Carlo (i.e., single level) estimator
@@
As stated before, it requires a computational complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})$ for the simple Monte Carlo to achieve the same level of MSE. From the view point of complexity, MLMC beats the simple Monte Carlo (i.e., the single level Monte Carlo) for all cases of $r$ and achieves the optimal complexity of $\mathcal{O}(\epsilon^{-2})$ when $r>\gamma$.
Nested simulation combined with the MLMC method has been studied for other applications [17, 18, 21, 29]. We next show how to use MLMC to estimate the gradient $\nabla_{\phi}\mathcal{L}(\phi)=\mathbb{E}\left[\rho_{\phi}\right]$, which is a nested expectation. To fit the MLMC setting, we can take $P=\rho_{\phi}$ and $P_{\ell}=\rho_{\phi,M_{\ell}}$, where $M_{\ell}=2^{\ell}M_{0}=2M_{\ell-1}$. It is clear that $P_{\ell}$ converges $P$ as $\ell\to\infty$. The core of MLMC is to construct coupling estimators $\Delta P_{\ell}$ with a variance decaying as $\mathcal{O}(2^{-r\ell})$. To gain a large $r$, we use the antithetic coupling construction as in [8, 18, 19, 21], which is better than the naive coupling $\Delta P_{\ell}=\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}$. The fundamental concept of antithetic coupling involves the selection of two non-overlapping subsets, each with a size of $M_{\ell-1}$ from the $M_{\ell}$ inner samples $\{\theta_{j}^{\prime}\}_{j=1}^{M_{\ell}}$ utilized for computing $\rho_{\phi,M_{\ell}}$ given in (14). This results in two independent realizations of $\rho_{\phi,M_{\ell-1}}$, identified as $\rho^{(a)}_{\phi,M_{\ell-1}}$ and $\rho^{(b)}_{\phi,M_{\ell-1}}$. Specifically,
@@
The notations $\rho^{(b)}_{\phi,M_{\ell-1}}$, $\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)$, and $g^{(b)}_{\phi,M_{\ell-1}}(x)$ are defined in the similar way by using $\theta^{\prime}_{j}$, $j=M_{\ell-1}+1,\dots,M_{\ell}$ instead. Define
@@
We have $\gamma=1$ for $\Delta\rho_{\phi,\ell}$ and $\Delta\psi_{\phi,\ell}$. We next examine the values of $\alpha$ and $r$. To this end, the next two theorems provide upper bounds for the first and second moments of $|\Delta\psi_{\phi,\ell}|$ and $\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ as $\ell\to\infty$.
@@
This proof follows an argument similar to Theorem 2 in [21], which particularly showed that $\mathbb{E}\left[|\Delta\psi_{\phi,\ell}|\right]=\mathcal{O}(2^{-\alpha_{1}\ell})$ with $\alpha_{1}=\min(p(q-1)/(2q),1)$, Since $(p-2)(q-2)\geq 4$, then we actually have $\alpha_{1}=1$. ∎
@@
we have the bias rate $\alpha=1$ for both cases. According to (19), the resulting MLMC estimator (18) yields the optimal complexity $\mathcal{O}(\epsilon^{-2})$. We should note that the nested APT estimator has a complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})=\mathcal{O}(\epsilon^{-3})$. Although the MLMC estimator (18) gains lower complexity than the nested APT estimator, it is biased and depends on the prespecified error threshold $\epsilon$. Moreover, the sample sizes $N_{\ell}$ for each level and the finest level $L$ need to be tuned by the threshold $\epsilon$. For practical training problems, it may be difficult to choose the threshold $\epsilon$. Since MLMC achieves the optimal regime $r>\gamma$, one can design unbiased MLMC estimators for eliminating the bias of the MLMC estimator (18). The unbiased MLMC also admits the optimal complexity $\mathcal{O}(\epsilon^{-2})$ without choosing the sample sizes $N_{\ell}$.

### v3

@@ Abstract
There is a growing interest in studying sequential neural posterior estimation (SNPE) techniques due to their advantages for simulation-based models with intractable likelihoods. The methods aim to learn the posterior from adaptively proposed simulations using neural network-based conditional density estimators. As an SNPE technique, the automatic posterior transformation (APT) method proposed by Greenberg et al. (2019) performs well and scales to high-dimensional data. However, the APT method requires computing the expectation of the logarithm of an intractable normalizing constant, i.e., a nested expectation. Although atomic proposals were used to render an analytical normalizing constant, it remains challenging to analyze the convergence of learning. In this paper, we reformulate APT as a nested estimation problem. Building on this, we construct several multilevel Monte Carlo (MLMC) estimators for the loss function and its gradients to accommodate different scenarios, including two unbiased estimators, and a biased estimator that trades a small bias for reduced variance and controlled runtime and memory usage. We also provide convergence results of stochastic gradient descent to quantify the interaction of the bias and variance of the gradient estimator. Numerical experiments for approximating complex posteriors with multimodality in moderate dimensions are provided to examine the effectiveness of the proposed methods.
@@ 1 Introduction
Simulator-based models are widely used across scientific disciplines, including neuroscience [46], physics [7, 23], biology [28, 38, 45], and inverse graphics [52]. These models serve as crucial tools for describing and understanding the data-generating process. However, when applying traditional Bayesian inference to simulator-based models, challenges arise, including an intractable likelihood function $p(x|\theta)$ and the computational expense associated with running the simulator.
@@
To address these challenges, a series of likelihood-free Bayesian computation (LFBC) methods have been developed. These methods include approximate Bayesian computation (ABC) [5, 44], synthetic likelihoods (SL) [50, 58], Bayesian optimization [26], likelihood-free inference by ratio estimation [56], and pseudo-marginal methods [2, 3]. A comprehensive summary and review of these methods can be found in [10], and they have all been benchmarked in [31, 40].
@@
Posterior density estimation approaches approximate the posterior of interest $p(\theta|x_{o})$ with a family of density estimators $q_{\phi}(\theta)$, where $\phi$ is the parameter to be tuned. Optimization-based approaches are widely used in these methods, and the Kullback-Leibler (KL) divergence between $p(\theta|x)$ and $q_{\phi}(\theta)$, which measures the difference between two densities, is commonly chosen as the loss function. Variational Bayes (VB), as an effective optimization-based method for approximating the posterior distribution of a Bayesian problem, is widely used. In the likelihood-free context, [57] developed a new VB method with an intractable likelihood, while [29] proposed an unbiased VB method based on nested MLMC. However, these methods tend to fail in cases where simulations are expensive, as the nested estimation in these methods requires additional simulation procedures.
@@
For this reason, there has been a growing interest recently in employing neural networks to represent probability densities, particularly normalizing flows [37]. When the inference problem focuses solely on the observation $x_{o}$, the data efficiency can be improved by using the sequential training schemes in the sequential neural posterior estimation (SNPE) methods [24, 41, 47]. In this approach, model parameters are drawn from a proposal distribution that is more informative about $x_{o}$ than the prior distribution. However, SNPE requires modifying the loss function relative to neural posterior estimation (NPE) to ensure that the neural network approximates the true posterior $p(\theta|x)$. Many approaches have been developed to address this issue.
Among SNPE methods, automatic posterior transformation (APT) requires computing the expectation of a logarithm. This logarithm is of an intractable normalizing factor, and crucially, the factor itself is also an expectation, resulting in a nested expectation overall. Greenberg et al. [24] then used atomic proposals to render an analytical normalizing constant. However, it remains challenging to perform convergence analysis using the existing analysis techniques [6]. To address this limitation, we reformulate APT as a nested estimation problem. We then develop several multilevel Monte Carlo (MLMC) [16] nested estimators for the loss function and its gradients, which enjoy lower computational complexity than the single-level nested estimator. In stochastic gradient training, the usual unbiased MLMC estimator [51] may suffer from high gradient variance and large GPU memory footprints. Motivated by [32], we study two MLMC variants: (i) an unbiased scheme that reduces variance at additional computational cost, and (ii) a truncated scheme that introduces a small, controlled bias to reduce variance while adhering to compute and memory budgets. We establish several theorems that investigate the order of bias, variance, and the average cost of the MLMC estimators. Finally, we establish convergence guarantees for stochastic gradient descent (SGD) and quantify the effects of the bias and variance of the gradient estimator.
@@
The remainder of this paper is organized as follows. In Section 2, we propose the nested APT method and analyze its computational complexity. In Section 3, we introduce the basics of MLMC methods, including their formulation and theoretical analysis of the random sequence utilized in MLMC methods. Equipped with these tools, we develop both unbiased MLMC methods and truncated MLMC methods and analyze the order of variance for their losses and gradients, as well as the average cost. In Section 4, we provide convergence analysis for nested APT and truncated MLMC methods in the case of SGD. In Section 5, we conduct a series of numerical experiments on benchmark tasks. Finally, we conclude this paper with some remarks in Section 6. Lengthy proofs and additional results are deferred to the appendices.
@@ 2.1 Problem formulation
Assume that the prior distribution has density $p(\theta)$ with respect to the Lebesgue measure, where $\theta\in\Theta$ denotes the model parameter of interest. We focus on the finite-dimensional parameter space $\Theta$. Given an observed sample $x_{o}$, our objective is to perform inference on the posterior $p(\theta|x_{o})\propto p(\theta)p(x_{o}|\theta)$. However, in many cases, the likelihood $p(x|\theta)$ either lacks a closed form or is difficult to evaluate directly. Instead, it can be expressed via a simulator so that we can sample $x$ from $p(x|\theta)$ for a fixed model parameter $\theta$.
@@
which is nonnegative by Jensen’s inequality and attains its minimum when $q(\theta)$ agrees with $p(\theta)$, making it suitable as a loss function. When the likelihood $p(x|\theta)$ is tractable, one can directly approximate $p(\theta|x_{o})$ by minimizing the KL divergence between the target distribution and the proposed estimator $q(\theta)$ within a certain family of distributions. In the likelihood-free context, this can be viewed as a problem of conditional density estimation. Within this framework, a conditional density estimator $q_{F(x,\phi)}(\theta)$ based on a neural network [47, 24] is utilized to approximate $p(\theta|x)$ over the admissible set of tuning parameters $\phi\in\Phi$, where $\Phi\subset\mathbb{R}^{d}$ is the space of the neural network parameters. To this end, we focus on minimizing the following average KL divergence under the marginal distribution with density $p(x)=\int p(\theta)p(x|\theta)\mathrm{d}\theta$
@@
is used as the loss function. Since $L_{1}(\phi)$ is intractable, we approximate it by the empirical estimator
@@
Since we aim to conduct conditional density estimation at $x_{o}$, it is essential to utilize a proposal $\tilde{p}(\theta)$ that is more informative regarding $x_{o}$ than the prior $p(\theta)$. After initializing $\tilde{p}(\theta)$ as $p(\theta)$, we then want the approximation of $p(\theta|x_{o})$ to serve as a good candidate for the proposal in subsequent simulations. This conditional density estimation with an adaptively chosen proposal is called sequential neural posterior estimation (SNPE). However, after replacing $p(\theta,x)$ with $\tilde{p}(\theta,x)=\tilde{p}(\theta)p(x|\theta)$ in (1), it is observed that $q_{F(x,\phi)}(\theta)$ approximates the proposal posterior
@@
where $\tilde{p}(x)=\int\tilde{p}(\theta)p(x|\theta)\mathrm{d}\theta$. Hence, we adjust the loss function so that $q_{F(x,\phi)}(\theta)$ approximates the true posterior $p(\theta|x)$.
@@
Since the integral $Z(x,\phi)$ is usually intractable in practice, APT proposes the use of ‘atomic’ proposals. APT with such proposals is known as atomic APT. Specifically, they assume that $\tilde{p}(\theta)$ is a discrete uniform distribution $\mathcal{U}\{\theta_{1},\dots,\theta_{M}\}$, where $\theta_{i}$ are sampled from a distribution. The uniform setting of the proposal distribution enables the analytical computation of $Z(x,\phi)$. With this method, (3) and (5) can be reformulated as
@@
However, to the best of our knowledge, the use of atomic proposals makes it challenging to analyze their convergence behavior. Therefore, we are unable to explain the unexpected low performance in some tasks [12] with the existing convergence results [6], both empirically and theoretically. As an alternative approach to estimating $Z(x,\phi)$, which enjoys a comprehensive theoretical framework and comparable performance, is studied in the next section.
@@ 2.2 Nested estimation
is selected as the loss function, and the last term $\mathbb{E}_{\tilde{p}(\theta)}\left[\log\tilde{p}(\theta)\right]$ in (7) is dropped for it is independent of $\phi$. We employ stochastic gradient methods to optimize the loss function in this paper, and therefore $\nabla_{\phi}\mathcal{L}(\phi)$ is of interest. By interchanging expectation $\mathbb{E}[\cdot]$ and gradient operator $\nabla_{\phi}$, the gradient of (8) is given by
@@
Notice that (9) is similar to the gradient examined in the optimization procedure of Bayesian experimental design [9, 15, 33, 36]. Given any query point $(\theta,x)\sim\tilde{p}(\theta,x)$, we denote the corresponding queries of the loss and gradient as
@@
Given the intractability of the normalizing constant $Z(x,\phi)=\mathbb{E}_{\tilde{p}(\theta^{\prime})}\left[g_{\phi}(x,\theta^{\prime})\right]$, it remains an obstacle to deriving an estimator for the loss function. A simple choice is leveraging its empirical estimator based on $M$ samples, which is given by
@@
$\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ $=\frac{1}{N}\sum_{i=1}^{N}\psi_{\phi,M}^{(i)}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{1}{M}\sum_{j=1}^{M}\frac{g_{\phi}(x_{i},\theta^{\prime}_{ij})}{g_{\phi}(x_{i},\theta_{i})},$ (15)
@@
where $\theta_{ij}^{\prime}\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta^{\prime})$ constitute the inner sample of size $M$ and $(\theta_{i},x_{i})\stackrel{{\scriptstyle iid}}{{\sim}}\tilde{p}(\theta,x)$ constitute the outer sample of size $N$, and $\psi^{(i)}_{\phi,M},\ \rho^{(i)}_{\phi,M}$ are iid copies of $\psi_{\phi,M}$, $\rho_{\phi,M}$ respectively. Due to the nonlinearity of the logarithm, (15) is a biased estimator. Using similar arguments to [53], it is straightforward to show that the nested estimator $\hat{{\mathcal{L}}}^{\mathrm{Ne}}(\phi)$ has a variance of $\mathcal{O}(1/N)$ and a bias of $\mathcal{O}(1/M)$. For a given mean squared error (MSE) $\epsilon$, choosing $N=\mathcal{O}(\epsilon^{-2})$ and $M=\mathcal{O}(\epsilon^{-1})$ yields a computational complexity of $\mathcal{O}(NM)=\mathcal{O}(\epsilon^{-3})$.
Plain nested estimators (13)–(16) are also known as single-level nested estimators compared to MLMC. Empirical results in Appendix C show that nested APT with the single-level nested estimator performs comparably to atomic APT. Casting the task as a nested estimation problem facilitates numerical analysis and further enables us to leverage MLMC techniques to enhance the single-level nested estimator.
@@ 3.1 Basic idea of MLMC
We first introduce the idea of MLMC in a generic setting following [16]. Consider the problem of estimating $\mathbb{E}[P]$, where the sampling of $P$ is costly or even infeasible. Suppose that $P$ can be approximated by tractable $P_{\ell}$ with increasing accuracy, but also increasing cost, as $\ell\to\infty$. As a result, the expectation $\mathbb{E}[P]$ can be estimated by the single-level Monte Carlo estimator,
@@
As stated before, it requires a computational complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})$ for the single-level Monte Carlo to achieve the same level of MSE. From the viewpoint of complexity, MLMC improves upon the single-level Monte Carlo for all cases of $r$ and achieves the optimal complexity of $\mathcal{O}(\epsilon^{-2})$ when $r>\gamma$.
Nested simulation combined with the MLMC method has been studied for other applications [17, 18, 21, 29]. We next show how to use MLMC to estimate the gradient $\nabla_{\phi}\mathcal{L}(\phi)=\mathbb{E}\left[\rho_{\phi}\right]$, which is a nested expectation. To fit the MLMC setting, we can take $P=\rho_{\phi}$ and $P_{\ell}=\rho_{\phi,M_{\ell}}$, where $M_{\ell}=2^{\ell}M_{0}=2M_{\ell-1}$. Note that the the base 2 is specifically chosen here solely for the convenience of subsequent analysis [22]. Since $P_{\ell}$ converges to $P$ as $\ell\to\infty$, the core of MLMC is to construct coupling estimators $\Delta P_{\ell}$ with a variance decaying as $\mathcal{O}(2^{-r\ell})$. To gain a large $r$, we use the antithetic coupling construction as in [8, 18, 19, 21], which is better than the naive coupling $\Delta P_{\ell}=\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}$. The fundamental concept of antithetic coupling involves the selection of two non-overlapping subsets, each of size $M_{\ell-1}$ from the $M_{\ell}$ inner samples $\{\theta_{j}^{\prime}\}_{j=1}^{M_{\ell}}$ utilized for computing $\rho_{\phi,M_{\ell}}$ given in (14). This results in two independent realizations of $\rho_{\phi,M_{\ell-1}}$, identified as $\rho^{(a)}_{\phi,M_{\ell-1}}$ and $\rho^{(b)}_{\phi,M_{\ell-1}}$. Specifically,
@@
The notations $\rho^{(b)}_{\phi,M_{\ell-1}}$, $\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)$, and $g^{(b)}_{\phi,M_{\ell-1}}(x)$ are defined in a similar way by using $\theta^{\prime}_{j}$, $j=M_{\ell-1}+1,\dots,M_{\ell}$ instead. Define
@@
We have $\gamma=1$ for $\Delta\rho_{\phi,\ell}$ and $\Delta\psi_{\phi,\ell}$, it remains to examine the values of $\alpha$ and $r$. To this end, the next two theorems provide upper bounds for the first and second moments of $|\Delta\psi_{\phi,\ell}|$ and $\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ as $\ell\to\infty$.
@@
This proof follows an argument similar to Theorem 2 in [21], which particularly showed that $\mathbb{E}\left[|\Delta\psi_{\phi,\ell}|\right]=\mathcal{O}(2^{-\alpha_{1}\ell})$ with $\alpha_{1}=\min(p(q-1)/(2q),1)$, Since $(p-2)(q-2)\geq 4$, we actually have $\alpha_{1}=1$. ∎
@@
we have the bias rate $\alpha=1$ for both cases. According to (19), the resulting MLMC estimator (18) yields the optimal complexity $\mathcal{O}(\epsilon^{-2})$. Note that the nested APT estimator has a complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})=\mathcal{O}(\epsilon^{-3})$. Although the MLMC estimator (18) achieves lower complexity than the nested APT estimator, it is biased and depends on the prespecified error threshold $\epsilon$. Moreover, the sample sizes $N_{\ell}$ for each level and the finest level $L$ need to be tuned by the threshold $\epsilon$. For practical training problems, it may be difficult to choose the threshold $\epsilon$. Since MLMC achieves the optimal regime $r>\gamma$, one can design unbiased MLMC estimators for eliminating the bias of the MLMC estimator (18). The unbiased MLMC also admits the optimal complexity $\mathcal{O}(\epsilon^{-2})$ without choosing the sample sizes $N_{\ell}$.